Tech Staff Meeting Notes May 4, 2015

The top priority is always repair of down stations.
NM problem stations
ET problem stations

Duty Table updated June 2014


Priorities


News & Highlights

The fish fry will on May 8 and cost $6.

We're getting 14 s-13's from South Carolina.

Chris is working on a noise reduction scheme for the maclab.

It is time to begin annual preventive maintenance for each station.

Jim asked about a heli_ew process for status channels. Mitch or James will add it to shakeworm. Its probably best for this heli_ew to get data directly from the nodes rather than set up and additional waveserver and ship the continous data to CERI.

Philip Crotwell is visiting May 7-8 and will bring S-13's that can be used to replace the NC L4's.


NMSZ

freewave serial reset installed at 9 of 11 broadbands.

We have 11 episensors cannabalized from the K2's. We'll replace the RT147's at all 11 NM broadband stations then cull the RT147's for working sensors. We'll use

Philip Crotwell is visiting May 7-8 and will bring S-13's that can be used to replace the NC L4's. those working sensors to replace any remaining malfuctioning RT147's and send the rest back to the Depot.

Mitch will contact Menlo to find out when they need their portable back. It looks like we successfully turned off the seismicity by installing a seismograph.

NMDown


Most images taken from the 2.9 event. /gaia/smeagol/dat2/Working/NM/201505/20150501_234120_MAN

Down
NMEM
RDST
LCAR.HH*
LCAR.HN*
WLAR.HN*
WLAR.HH*

Misc Problem
HENM.HNE - spikes
HENM.HNZ - spikes
PARM.HN* - see image
LNXT.HNZ.00 - timing- coming in too early -see image
BOAR and LVAR -gaps since one repeats through the other
HDBT.HNE- ? see image
SEAR.HNE - ? see image

Two ISIS3 upgrades remain: ARPT, and the PLAR replacement.

The Basalt is ready for installation at CUET. Done.

RDST has a very large pole right next to the vault that will corrupt any data during strong ground motion. It needs to get moved; Dave has a couple options on the Rhodes campus (the name will change when its moved). Hopefully rather than reinstall the Guralps in the new location, we can install a Basalt instead.

NODE Tasks

The win7 replacement for mudworm will be installed this week. Its a 32 bit installation so we don't need to upgrade memory. So they suffer from the Unix Millenium Bug and their epoch times will reset to 1901 in 2038: http://en.wikipedia.org/wiki/Year_2038_problem.

DBB:

Initial test of DBB through Wilder went well. But there's too much torque in the tower and we lose connection in the wind. Jim got permission to add a stiffener to the tower.

MKTA:

LNXT:

NMAD:

The waveserver continues to have occasional very gappy periods most likely due to missed messages. Steve will investigate.

AZTN/HDBR: Shahram submitted a proposal to TDOT to replace all the sensors with Episensors (including new cabinets, etc). We asked USGS to provide new radios so we can have a continuous buffer of all channels at AZTN. We'll then allow access to NSMP so they can extract event data.

CERI:

Need to review and update the power systems in 218.


ETSZ

John is helping Gordana Vlahovic get NCCU back online. She has a computer that needs to be configured. Its best to use earthworm exports rather than scream to scream. Steve will help get this done.

CCNC is not going through fgtn and instead repeats through ASTN. The 950MHz shot would work so it could be converted to digital.

We're still repeating CPRT through GTTN. The plan was to put a cell modem at CPRT but John found a potential internet host in the local Library.

TZTN is repaired. John has a plan to propose to USGS to harden against critters.

PMTN was vandalized by copper thieves. John spoke to the Sheriff who advised that once they do it once, its likely to happen again. John will look for alternative locations. We might still be able to use PMTN as a repeater/internet host. SMNC has a bad horizontal. Replace L4 with S-13.

ETDown


Images taken from here: /gaia/smeagol/dat2/Working/ET/201505/20150502_120210_MAN

Down
PMTN.HN*
PMTN.HH*
VHTN.EHE
CCRT.EHE

Misc Problem
BCRT.HN*-spikes
CPRT.EHE- drifting
CPRT.HN* -spikes
ASTN.HNN - see image

STAR Node We may have a bad intrusion alarm. Should put a better lock on the cabinet.

Lightning at CCRT has consumed two refteks, an S-13, and several other other pieces of hardware despite having the best grounding and surge supression of any station in the network. We're going to cut our losses and downgrade to ISIS4. The reftek and episensor will find a new home, possibly at CCNC.

The unused pair of ubiquiti radios will be installed at Star Mountain. All CERI data will then go over the Ubiquiti link and the EMA can use the Aironet link to transmit the camera data in real-time continuous.

BRBC Node

Morgantown node hardware is ready to go (needs an earthworm) and waiting on completion of the ice bridge and fiber.

Internet, accessibility, and high noise inducing RF requires we move the MGNC central receive. John presented several possibilities and investigated both the PSN digitizer and the Menlo Park system. We determined that eventually we'll need to convert all nodes to the Menlo Park system so that we can obsolesce XP. But for now, we need to reconfigure North Carolina as in this figure. Better figure. This includes:
Decommission RBNC DONE
Upgrade WSNC to broadband or S-13/Reftek and cell modem. Done.
Rotate CCNC to repeat through fgtn and add to MTTN node. Done.
Establish new node at High Peak.
Transmit MGNC to High Peak. Done
Repeat TRYN through MGNC. Too much RF, will test SMNC instead
Transmit GFM to High Peak.
Transmit TVNC to High Peak.
Transmit SMNC to High Peak. Done

Node hardware at BRBC is pulled. Waiting on tower work at Morgantown. John volunteered to remove the old hardware and install the new, but the tower should be repainted.

MTTN Node

CPRT is waiting on county for permission to use Library internet; will probably be ok.

GMG needs a new vertical S-13.

PMTN needs a fence?


GPS

Jim provided an update to the spreadsheet detailing the antenna upgrades and ethernet isolation progress at each station. All but two are complete (CJTR and HCES).

CJTR is still offline. Waiting for contractors to complete work at the site.


AG Network

See comments in the News and Highlights section on power upgrades.

LCAR got water in again. The trillium appears to be ok but the RT147 is not. The source of the leak is the conduit into the vault.

WLAR got a new router but still does not connect to the internet. Needs a site visit.


FDL

The giddings has a new cylinder and two new tires.

The committee provided the completed report that was discussed at the February faculty meeting. Chuck would like to see a well equipped, well orgranized and well used FDL. All that takes is time and money of which we have neither. Outside of JerMing's class and Steve Horton's portables, the equipment isn't used much. The FDL wiki is here: https://umwiki.memphis.edu/display/FDL/CERI+FDL


NSMP

NSMP Station List.


MISCELLANEOUS

The seisnet group on umdrive is at http://umdrive.memphis.edu/mwithers/seisnet You will need your uuid and password.